# QMill: Representative Quantum Data Generation for Quantum Machine Learning Utility

### v2

@@ 5.3 QMill’s Evaluation Metrics
6 QMill’s Evaluation and Analysis
6.1 QMill’s Ability to Capture Distributions
We evaluate QMill across multiple CE distributions, observing varied performance depending on the target shape. Fig. 5 presents the best-performing ansatz for each case. For the uniform distribution, ansatz A1 achieves a TVD of $\approx 0.18$, reflecting reasonable spread coverage. In the Gaussian case, ansatz A4 performs best, achieving a TVD below $0.1$ (Fig. 6(a)). The trained distribution accurately captures both the central peak and the bell-shaped spread, closely matching the target. Ansatz A1 also performs well on the left-skewed Weibull, effectively modeling the sharp peak and gradual decline. The right-skewed Weibull, however, proves more challenging: although ansatz A3 reduces the TVD to 0.5, the improvement over the initial state is modest. This distribution intentionally concentrates CE at unrealistically high values to stress test QMill’s limits. Despite these extremes, QMill achieves reasonably low TVD across all cases, demonstrating robustness even under adversarial conditions.
Figure 6: QMill’s TVD performance on (a) arbitrary distributions used for stress testing its impact and (b) real datasets (lower is better).
6.2 QMill’s Ability to Emulate Real Datasets
QMill shows strong performance when emulating CE distributions from real-world classical and quantum datasets. Across all evaluated datasets, the trained distributions align closely with targets, with high-fidelity matches observed in most cases. For MNIST, ansatz A3 achieves a TVD << 0.1, significantly outperforming A1 and A2 and accurately reproducing the characteristic bell-shaped CE profile (Fig. 5, Fig. 6(b)). Similar performance is observed for FashionMNIST and CIFAR, with QMill consistently narrowing the initial CE spread to better match the target structure.
Figure 7: QMill’s TVD with quantum sensors.
On quantum datasets, QMill performs especially well. For the quantum chemistry dataset, all ansatzes yield TVD values below 0.2 despite the narrow CE band, and results for soil moisture and dark matter datasets similarly show close alignment (Fig. 5, Fig. 7). While later evaluations show some ansatzes outperform others overall, these results highlight that different architectures excel on specific distributions. For example, A3 is best for MNIST, A2 performs well on soil and DM sensor signals, A1 is optimal for Left Weibull, and A4 captures the chemistry dataset most effectively. This underscores the utility of maintaining a diverse ansatz library tailored to different CE profiles.
@@ 6.2 QMill’s Ability to Emulate Real Datasets
6.3 Diversity of Samples Generated by QMill
We assess the diversity of generated states using SWAP tests between state pairs within similar CE ranges across all four circuit architectures. As shown in Fig. 8(a), each point represents the average SWAP test value for a given CE range, with point size indicating the number of state pairs tested (the larger the circle, the more the samples). Most values lie between 0.5 and 0.6, suggesting that generated states are largely distinct, even within the same CE bin. We observe slightly higher similarity in the low CE regime, where most samples are concentrated—an expected outcome, as high-CE states are harder to generate. In contrast, states in higher CE ranges consistently yield SWAP scores near 0.5, indicating strong sample-level diversity. This trend is consistent across all ansatzes, confirming that QMill reliably produces non-redundant states across the full CE spectrum.
6.4 QMill’s Performance Under Noise
@@ 6.5 Performance of Different Ansatzes
6.6 Deploying QMill for QML Classifiers
We now demonstrate QMill practical utility, specifically whether these synthetic CE datasets can effectively train QML models. We train a three-qubit QNN on QMill-generated CE feature sets under both ideal and noisy simulations, and benchmark its accuracy against a classical logistic-regression ceiling. We first create a dataset using the quantum soil sensor data we generated earlier by batching the CE values into 400 samples, with each sample containing 9 CE values, and then assigning the label 0 or 1 depending on whether the samples came from low vs. high moisture soil. We implement a three‐qubit classifier by encoding each of the input CE value features into an RY–RX–RZ feature map, then applying an ansatz with full entanglement between qubits using Qiskit’s RealAmplitudes parameterized circuit. The model then measures a single‐qubit Z observable on the first qubit and feeds the expectation value into a QNN. Training is performed with a dual‐annealing optimizer, and we evaluate performance with 5-fold cross-validation. Noise is modeled using IBM Sherbrooke’s error parameters. Table 1 summarizes accuracy, precision, recall, and F1 score for the noisy and ideal circuits, each normalized against a classical logistic‐regression baseline set to 100%. Notably, the noisy implementation falls within a few percentage points of its ideal counterpart, demonstrating that our three-qubit classifier retains nearly all of its predictive power even in the presence of realistic gate and readout errors. This tight correspondence confirms that, for Concentratable Entanglement–based features, the modest noise levels expected on near-term quantum hardware will not adversely affect QMill’s performance for QML applications.
@@ 7 Related Work
As QMill is the first-of-its-kind effort toward synthetic QML data generation, the related work is limited. Schatzki et al. (Schatzki et al., 2021) attempted to generate entangled datasets using quantum circuits trained to achieve a single target, concentratable entanglement value. However, this approach falls short as generated samples often deviate from the desired entanglement. Xu et al. (Xu et al., 2025) employed supervised QML and CE lower bound metric to generate mixed-state datasets designed for entanglement classification around a target value, which is orthogonal to our approach of generating target CE distribution datasets. Zhang et al. (Zhang et al., 2025) uses a denoising model to synthesize class-specific GHZ/W-like states; unlike QMill, this does not control CE distributions across datasets nor enforce sample diversity.
Other approaches include domain-specific methods, such as Quantum Generative Adversarial Networks (QGANs) for detecting product states (Steck & Behrman, 2024), and quantum transfer learning on small, high-dimensional datasets for remote sensing (Otgonbaatar et al., 2023). While innovative, these methods don’t generalize for QML tasks requiring flexible entanglement distributions. Yu et al. (Yu et al., 2023) proposed generating optimal datasets for learning unitary transformations, yet the approach remains constrained to classical applications. Sim et al. (Sim et al., 2019) explored the expressibility of parameterized quantum circuits, providing insight into ansatz selection, but in our work, we observe that higher expressibility does not necessarily correlate with better CE matching. This limitation necessitated the design of a customized ansatz in QMill to better align with targeted CE distributions, enabling more effective synthetic data generation across a range of entanglements.
@@ 8 Conclusion
We introduced QMill, a quantum data generation framework that produces diverse datasets with distributions of concentratable entanglement values, supporting robust QML model development. By leveraging customizable ansatz and efficient, low-depth circuits with SWAP tests, QMill enables scalable, high-quality synthetic data generation with a diverse set of samples, validated across multiple classical and quantum datasets. QMill thus addresses a critical need in QML, providing an essential quantum data generation framework for advancing QML training and evaluation to attain quantum utility and speedup in practice.
@@ 9 Acknowledgement
This work was supported by Rice University, the Rice University George R. Brown School of Engineering and Computing, and the Rice University Department of Computer Science. This work was supported by the DOE Quantum Testbed Finder Award DE-SC0024301, the Ken Kennedy Institute, and Rice Quantum Initiative, which is part of the Smalley-Curl Institute. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.

### v3

@@ 5.3 QuFoundry’s Evaluation Metrics
6 QuFoundry’s Evaluation and Analysis
6.1 QuFoundry’s Ability to Capture Distributions
We evaluate QuFoundry across multiple CE distributions, observing varied performance depending on the target shape. Fig. 5 presents the best-performing ansatz for each case. For the uniform distribution, ansatz A1 achieves a TVD of $\approx 0.18$, reflecting reasonable spread coverage. In the Gaussian case, ansatz A4 performs best, achieving a TVD below $0.1$ (Fig. 6(a)). The trained distribution accurately captures both the central peak and the bell-shaped spread, closely matching the target. Ansatz A1 also performs well on the left-skewed Weibull, effectively modeling the sharp peak and gradual decline. The right-skewed Weibull, however, proves more challenging: although ansatz A3 reduces the TVD to 0.5, the improvement over the initial state is modest. This distribution intentionally concentrates CE at unrealistically high values to stress test QuFoundry’s limits. Despite these extremes, QuFoundry achieves reasonably low TVD across all cases, demonstrating robustness even under adversarial conditions.
Figure 6: QuFoundry’s TVD performance on (a) arbitrary distributions used for stress testing its impact and (b) real datasets (lower is better).
6.2 QuFoundry’s Ability to Emulate Real Datasets
QuFoundry shows strong performance when emulating CE distributions from real-world classical and quantum datasets. Across all evaluated datasets, the trained distributions align closely with targets, with high-fidelity matches observed in most cases. For MNIST, ansatz A3 achieves a TVD << 0.1, significantly outperforming A1 and A2 and accurately reproducing the characteristic bell-shaped CE profile (Fig. 5, Fig. 6(b)). Similar performance is observed for FashionMNIST and CIFAR, with QuFoundry consistently narrowing the initial CE spread to better match the target structure.
Figure 7: QuFoundry’s TVD with quantum sensors.
On quantum datasets, QuFoundry performs especially well. For the quantum chemistry dataset, all ansatzes yield TVD values below 0.2 despite the narrow CE band, and results for soil moisture and dark matter datasets similarly show close alignment (Fig. 5, Fig. 7). While later evaluations show some ansatzes outperform others overall, these results highlight that different architectures excel on specific distributions. For example, A3 is best for MNIST, A2 performs well on soil and DM sensor signals, A1 is optimal for Left Weibull, and A4 captures the chemistry dataset most effectively. This underscores the utility of maintaining a diverse ansatz library tailored to different CE profiles.
@@ 6.2 QuFoundry’s Ability to Emulate Real Datasets
6.3 Diversity of Samples Generated by QuFoundry
We assess the diversity of generated states using SWAP tests between state pairs within similar CE ranges across all four circuit architectures. As shown in Fig. 8(a), each point represents the average SWAP test value for a given CE range, with point size indicating the number of state pairs tested (the larger the circle, the more the samples). Most values lie between 0.5 and 0.6, suggesting that generated states are largely distinct, even within the same CE bin. We observe slightly higher similarity in the low CE regime, where most samples are concentrated—an expected outcome, as high-CE states are harder to generate. In contrast, states in higher CE ranges consistently yield SWAP scores near 0.5, indicating strong sample-level diversity. This trend is consistent across all ansatzes, confirming that QuFoundry reliably produces non-redundant states across the full CE spectrum.
6.4 QuFoundry’s Performance Under Noise
@@ 6.5 Performance of Different Ansatzes
6.6 Deploying QuFoundry for QML Classifiers
We now demonstrate QuFoundry practical utility, specifically whether these synthetic CE datasets can effectively train QML models. We train a three-qubit QNN on QuFoundry-generated CE feature sets under both ideal and noisy simulations, and benchmark its accuracy against a classical logistic-regression ceiling. We first create a dataset using the quantum soil sensor data we generated earlier by batching the CE values into 400 samples, with each sample containing 9 CE values, and then assigning the label 0 or 1 depending on whether the samples came from low vs. high moisture soil. We implement a three‐qubit classifier by encoding each of the input CE value features into an RY–RX–RZ feature map, then applying an ansatz with full entanglement between qubits using Qiskit’s RealAmplitudes parameterized circuit. The model then measures a single‐qubit Z observable on the first qubit and feeds the expectation value into a QNN. Training is performed with a dual‐annealing optimizer, and we evaluate performance with 5-fold cross-validation. Noise is modeled using IBM Sherbrooke’s error parameters.
Table 1 summarizes accuracy, precision, recall, and F1 score for the noisy and ideal circuits, each normalized against a classical logistic‐regression baseline set to 100%. Notably, the noisy implementation falls within a few percentage points of its ideal counterpart, demonstrating that our three-qubit classifier retains nearly all of its predictive power even in the presence of realistic gate and readout errors. This tight correspondence confirms that, for Concentratable Entanglement–based features, the modest noise levels expected on near-term quantum hardware will not adversely affect QuFoundry’s performance for QML applications.
@@ 7 Related Work
As QuFoundry is the first-of-its-kind effort toward synthetic QML data generation, the related work is limited. Schatzki et al. (Schatzki et al., 2021) attempted to generate entangled datasets using quantum circuits trained to achieve a single target, concentratable entanglement value. However, this approach falls short as generated samples often deviate from the desired entanglement. Xu et al. (Xu et al., 2025) employed supervised QML and CE lower bound metric to generate mixed-state datasets designed for entanglement classification around a target value, which is orthogonal to our approach of generating target CE distribution datasets. Zhang et al. (Zhang et al., 2025) uses a denoising model to synthesize class-specific GHZ/W-like states; unlike QMill, this does not control CE distributions across datasets nor enforce sample diversity.
Other approaches include domain-specific methods, such as Quantum Generative Adversarial Networks (QGANs) for detecting product states (Steck and Behrman, 2024), and quantum transfer learning on small, high-dimensional datasets for remote sensing (Otgonbaatar et al., 2023). While innovative, these methods don’t generalize for QML tasks requiring flexible entanglement distributions. Yu et al. (Yu et al., 2023) proposed generating optimal datasets for learning unitary transformations, yet the approach remains constrained to classical applications. Sim et al. (Sim et al., 2019) explored the expressibility of parameterized quantum circuits, providing insight into ansatz selection, but in our work, we observe that higher expressibility does not necessarily correlate with better CE matching. This limitation necessitated the design of a customized ansatz in QuFoundry to better align with targeted CE distributions, enabling more effective synthetic data generation across a range of entanglements.
@@ 8 Conclusion
We introduced QuFoundry, a quantum data generation framework that produces diverse datasets with distributions of concentratable entanglement values, supporting robust QML model development. By leveraging customizable ansatz and efficient, low-depth circuits with SWAP tests, QuFoundry enables scalable, high-quality synthetic data generation with a diverse set of samples, validated across multiple classical and quantum datasets. QuFoundry thus addresses a critical need in QML, providing an essential quantum data-generation framework to advance QML training and evaluation and enable quantum utility and speedups in practice.
@@ 9 Acknowledgement
This work was supported by Rice University, the Rice University George R. Brown School of Engineering and Computing, and the Rice University Department of Computer Science. This work was supported by the DOE Quantum Testbed Finder Award DE-SC0024301, the Ken Kennedy Institute, and Rice Quantum Initiative, which is part of the Smalley-Curl Institute. LLMs were used for text editing and coding in this work, and all outputs were reviewed by the authors. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum team.